OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation y = 4x + 5? a. 4 b. –4 c.-1/4 d. –5

OpenStudy (amoodarya):

m *m'=-1 when slope m=4 so m'=?

OpenStudy (amoodarya):

can you give the answer?

OpenStudy (anonymous):

-4?

OpenStudy (amoodarya):

no (4)(-4)= ?-1

OpenStudy (amoodarya):

i give example y=3x+2 is perpendicular to y=-1/3x+6

OpenStudy (amoodarya):

slope of perpendicular line is ?

OpenStudy (amoodarya):

4*(?)=-1

OpenStudy (amoodarya):